### v4

# A Survey of Deep Learning for Geometry Problem Solving

Abstract
Geometry problem solving is a key area of mathematical reasoning, which is widely involved in many important fields such as education, mathematical ability assessment of artificial intelligence, and multimodal ability assessment. In recent years, the rapid development of deep learning technology, especially the rise of multimodal large language models, has triggered a widespread research boom. This paper provides a survey of the applications of deep learning in geometry problem solving, including (i) a comprehensive summary of the relevant tasks in geometry problem solving; (ii) a thorough review of related deep learning methods; (iii) a detailed analysis of evaluation metrics and methods; and (iv) a critical discussion of the current challenges and future directions that can be explored. Our goal is to provide a comprehensive and practical reference of deep learning for geometry problem solving to promote further developments in this field. We create a continuously updated list of papers on GitHub: https://github.com/majianz/dl4gps.
A Survey of Deep Learning for Geometry Problem Solving
Jianzhe Ma1    Wenxuan Wang1∗    Qin Jin1thanks:   Qin Jin and Wenxuan Wang are the corresponding authors. 1Renmin University of China {majianzhe, wangwenxuan, qjin}@ruc.edu.cn
1 Introduction
As a core component of mathematical reasoning, Geometry Problem Solving (GPS) has been closely related to the field of education and the evaluation of mathematical ability of AI for many years Narboux et al. (2018). Due to the close connection between geometry problems and geometric diagrams, it has naturally become a typical representative of multimodal mathematical problems. Solving geometry problems in the format of educational exams not only requires artificial intelligence systems to understand and interpret geometric diagrams, but also requires strong logical reasoning and numerical calculation capabilities, which makes it an ideal scenario for evaluating the perception and reasoning capabilities of deep learning models. In recent years, the rise of Multimodal Large Language Models (MLLMs) has further promoted the development of this field, fully demonstrating the great potential of deep learning models in handling complex geometry-related solving and visual understanding tasks. The number of papers on deep learning for geometry problem solving grows approximately from 1 in 2018 to 110 in 2024 and continues to increase in 2025 (see Figure 5).
Figure 1: Overview of the survey’s structure
Although there have been many surveys of deep learning methods and Large Language Models (LLMs) in the field of mathematical reasoning Lu et al. (2023); Sun et al. (2023); Ahn et al. (2024); Zhang et al. (2024j); Saraf et al. (2024); Yan et al. (2024); Kurisappan and Pandiya (2025), there is still a gap in the specialized field of geometry problem solving compared to other mathematical fields Zhang (2022); Li et al. (2024e). Recent geometry problem solving surveys are relatively limited, either focusing only on multimodal plane geometry problems Cho et al. (2025b) or failing to provide a comprehensive summary of deep learning datasets and methods Zhao et al. (2025b).
In this study, we started with classic papers in this field, used a single round of forward and backward snowballing, searched Google Scholar with the keyword “geometry”, and manually screened to ensure the relevance of the papers. In the end, we collected more than 310 academic papers that involved deep learning for geometry problem solving, and conducted a comprehensive and in-depth survey.
In the following sections, we will first summarize the tasks related to geometry problem solving in depth (§2). Then, we will comprehensively review the various methods used in the field of geometry problem solving (§3). After that, we perform a systematic analysis of the evaluation metrics and methods (§4). Finally, we will discuss the current challenges facing this field and look forward to future development directions (§5).
2 Geometry Problem Solving Tasks
In this section, we briefly introduce the tasks related to geometry problem solving. These tasks can be divided into fundamental tasks, core tasks, and composite tasks. Fundamental tasks involve the basic abilities required to solve geometry problems, core tasks are directly related to geometry problem solving, and composite tasks consider geometry problem solving as a component of other complex tasks. The taxonomy of tasks and datasets for geometry problem solving is presented in Figure 2, and a detailed summary of the datasets can be found in Table 1 and Table 2.
{forest}
[…]
Figure 2: Taxonomy of Tasks and Datasets for Geometry Problem Solving.
2.1 Fundamental Tasks
In order to solve geometry problems, a deep learning system must first have a variety of fundamental capabilities, including understanding geometric diagrams, semantic parsing of geometry problem texts, extraction of geometric relationships, and prediction of geometric knowledge.
Geometric Diagram Understanding. Geometric diagram understanding is committed to fully understanding the information in geometric diagrams. The subtasks it contains can be divided into multiple levels. First, identify and detect basic geometric elements (such as points, lines, angles, and polygons) and their basic attributes (such as quantity and size) Lu et al. (2015); Song et al. (2017, 2020). This process is called Geometric Element Recognition. Second, on the basis of the identification of geometric elements, further construct the structure and spatial relationship between elements Xia and Yu (2021); Huang et al. (2023), namely Geometric Structure Recognition. Some work refers to these two parts as Geometric Perception tasks Kamoi et al. (2024); Xing et al. (2024). Third, based on geometric perception ability, generate formal language for geometric diagrams Hao et al. (2022); Wei et al. (2024). This process is also called Geometric Diagram Parsing. Finally, some studies use natural language to provide an accurate description of geometric diagrams. These descriptions are either generated based on diagram parsing or directly generated from geometric diagrams Zhang and Moshfeghi (2024); Huang et al. (2025f), which is called Geometric Diagram Captioning.
Semantic Parsing for geometry problem texts. Semantic parsing is a key technology for converting problem text into machine-readable formal statements Matsuzaki et al. (2017), and is an important component of the early deep learning framework for geometry problem solving Joshi et al. (2018); Sun et al. (2019); Jayasinghe and Ranathunga (2020); Gonzalez et al. (2021). Geometry problem texts usually consist of multiple sentences and contain complex geometric knowledge information, and it is difficult to process cross-sentence references and information from the geometric domain Hopkins et al. (2017). There are also studies that combine geometric diagram parsing with semantic parsing, aiming to achieve joint parsing of text and images Boob et al. (2023); Zhou et al. (2024c).
Geometric Relation Extraction. Geometric relation extraction is a well-defined task that includes extracting geometric relationships from the question text Huang et al. (2022), or extracting high-confidence geometric relationships from text and diagrams, respectively Gan et al. (2017), and then storing them in a special format such as triples Zhou et al. (2022) and knowledge graphs Wang et al. (2025h). The model achieves a deep understanding of the problem by extracting geometric relationships in geometry problems rather than using natural language Gan et al. (2019b, a). Geometric Knowledge Prediction. Geometric knowledge prediction aims to evaluate the model’s geometric understanding ability by predicting the geometric principles Xu et al. (2025b) and theorems Lu et al. (2021) (i.e., geometric knowledge) required to solve geometry problems Ning et al. (2025). The model needs to predict the relevant geometric knowledge required to solve the problem based on the input question and apply it in the reasoning process Wu et al. (2024a).
2.2 Core Tasks
Geometry problem solving can be divided into two categories: geometry theorem proving and geometric numerical calculation Chen et al. (2022). On the premise of having the capabilities implied by the fundamental tasks, the model needs to solve geometry problems in the form of educational exams. See Figure 3 for an example.
Geometry Theorem Proving. Geometry theorem proving is a long-standing task in the field of artificial intelligence Gelernter et al. (1960); Kapur (1986). Its input is a geometry problem that needs to be proved, and its goal is to output a detailed derivation process of the proof, usually focusing on plane geometry.
Geometric Numerical Calculation. Geometric numerical calculation has gradually emerged with the introduction of new datasets in recent years Seo et al. (2015); Sachan et al. (2017). The input is a geometry problem involving the calculation of a certain geometric value (such as length or angle), and the desired output is a concise answer to the problem, without necessarily providing a complete reasoning process. Its question types can usually be divided into plane geometry, solid geometry, analytic geometry, etc.
Figure 3: An example of geometry theorem proving and geometric numerical calculation problem.
2.3 Composite Tasks
Recently, geometry problem solving has also often appeared as a subtask of comprehensive tasks, mainly used to explore the model’s ability in mathematical reasoning.
Mathematical Reasoning. Geometry is an important part of mathematics, and geometric diagrams are also a typical type of mathematical synthetic image. Therefore, geometry problems are often included in single-modal or multi-modal mathematical Hendrycks et al. (2021); Lu et al. (2024) and scientific Jiang et al. (2024) reasoning benchmarks to evaluate the performance of models in mathematical reasoning tasks.
3 Methods for Geometry Problem Solving
This section provides a comprehensive review of deep learning methods for geometry problem solving. First, we will introduce the relevant deep learning model architecture, and then systematically classify and summarize other deep learning methods from the training stage and the inference stage. The taxonomy of these methods is shown in Figure 4.
{forest}
[…]
Figure 4: Taxonomy of Deep Learning Methods for Geometry Problem Solving.
3.1 Architectures for Geometry Problem Solving
In geometry problem solving, the classic deep learning architecture is usually the Encoder-Decoder architecture Sutskever et al. (2014), which also includes MLLMs that have been widely used in recent years. In addition, many other architectures have also been applied, such as Generative Adversarial Networks (GANs) Goodfellow et al. (2014), Graph Neural Networks (GNNs) Scarselli et al. (2008), and Decoder-Only architectures. These architectures are outlined in more detail in Table 4 in the Appendix.
3.1.1 Encoder-Decoder Architecture
The encoder-decoder architecture can be divided into the following five key parts: text encoder, diagram encoder, multimodal fusion module, decoder, and optional knowledge module.
Text Encoder. Text encoder can convert the text content of the geometry problem into formalized statements or encode it into vectors, so that the deep learning system can understand and process the text information. In addition to rule-based methods Lu et al. (2021), early research works usually use Recurrent Neural Networks (RNNs) Elman (1990) to parse Joshi et al. (2018); Gonzalez et al. (2021) or encode Tsai et al. (2021); Chen et al. (2021) geometry problem texts. Common models include Long Short-Term Memory network (LSTM) Hochreiter and Schmidhuber (1997), Gated Recurrent Unit (GRU) Cho et al. (2014), and their bidirectional variants - BiLSTM and BiGRU. Some works use Transformer Vaswani et al. (2017) to encode text Zhang et al. (2023a); Ma et al. (2024a); Zhang et al. (2024f). In addition, there are also research works that use pre-trained language models for text encoding Jian et al. (2023b); Huang et al. (2022); Zhu et al. (2025), such as BERT Devlin et al. (2019) and T5 Raffel et al. (2020). The dual encoder structure of RoBERTa Liu et al. (2019) plus BiLSTM also shows good results Cao and Xiao (2022); Ning et al. (2023); Xiao et al. (2024a); Zhang et al. (2024a).
Diagram Encoder. Parsing geometric diagrams into formal statements or encoding them into vector information is of great significance for solving multimodal geometry problems. Early studies mainly use Convolutional Neural Networks (CNNs) LeCun et al. (1998) to encode geometric diagrams Zhang et al. (2023a, 2024f); Zhang and Moshfeghi (2024), and the network architectures involved include RetinaNet Lin et al. (2017) and its DenseNet Huang et al. (2017) variants Lu et al. (2021); Guo and Jian (2022); Jian et al. (2023a); Huang et al. (2022); Ma et al. (2024a), ResNet He et al. (2016) and its ConvNeXt Liu et al. (2022) variants Chen et al. (2021); Cao and Xiao (2022); Zhang et al. (2024a, d), and Fast R-CNN Girshick (2015); Jian et al. (2023b). Recent studies have widely adopted pre-trained diagram encoders, such as ViT Dosovitskiy et al. , ViTMAE He et al. (2022), CLIP-ViT Radford et al. (2021), SigLIP Zhai et al. (2023), and Swin-Transformer Liu et al. (2021), which are mainly used to build MLLMs. In addition, Iordan (2022), Zhang et al. (2022b) and Zhu et al. (2025) use LSTM, GNN and BLIP Li et al. (2022) to parse geometric diagrams respectively, while UniMath Liang et al. (2023) encodes diagrams through VQVAE Van Den Oord et al. (2017).
Some other studies use a CNN-Transformer architecture to combine the functions of a text encoder and a diagram encoder into a multimodal encoder Li et al. (2024h); Lin et al. (2025).
Multimodal Fusion Module. If the geometry problem is multimodal, the multimodal fusion module can fuse and align the text and diagram information extracted from the original geometry problem or encoder, and then pass the information to the decoder. Inspired by Yu et al. (2019), many studies introduce a co-attention module to comprehensively fuse and align text and image representations Chen et al. (2021); Cao and Xiao (2022); Ning et al. (2023); Pan et al. (2023); Ma et al. (2024a). Many MLLMs are also equipped with multimodal fusion modules to improve their multimodal understanding capabilities, for example, LLaVA-v1.5 Liu et al. (2024a) and MAmmoTH-VL Guo et al. (2024a) both use a two-layer MLP visual-language connector Shi et al. (2024); Li et al. (2024g); Xu et al. (2024b); Sharma et al. (2025); Ning et al. (2025); Jia et al. (2025); GLM-4V GLM et al. (2024) and Qwen2.5-VL Qwen et al. (2025) use MLP to map image representations to text space Yang et al. (2024b); Pan et al. (2025); Peng et al. (2025); and InternVL2 Chen et al. (2024d) uses the QLLaMA architecture Deng et al. (2024); Xu et al. (2025a). Some studies consider this module and the subsequent decoder as an overall encoder-decoder structure Jian et al. (2023b); Zhang et al. (2023a); Li et al. (2024h); Lin et al. (2025); Liang et al. (2023); Zhang and Moshfeghi (2024), using self-attention units, BiGRU and T5-Encoder.
Decoder. This module decodes the geometric knowledge and information to output the final answer to the question. Many studies use LSTM Chen et al. (2021); Cao and Xiao (2022); Ning et al. (2023); Pan et al. (2023); Xiao et al. (2024a); Zhang et al. (2024a); Ma et al. (2024a) or GRU Tsai et al. (2021); Jian et al. (2023b); Zhang et al. (2023a); Li et al. (2024h); Zhang et al. (2024d, f) as decoders in deep learning systems, which may also integrate attention mechanisms. Other studies use pre-trained language models as decoders. For example, Liang et al. (2023) and Zhang and Moshfeghi (2024) used T5-Decoder, Pan et al. (2024) chooses BERT, Peng et al. (2024) uses DeepSeekMath-RL Shao et al. (2024), and Zhuang et al. (2025); Shengyuan and Xiuqin (2024); Zhang et al. (2024c) use the Qwen Bai et al. (2023) series model. In addition, Zhang et al. (2025c) uses MAmmoTH2 Yue et al. (2024b), Zhang et al. (2025e) chooses Yi-1.5 Young et al. (2024), and Cho et al. (2025a) uses Llama 3 Grattafiori et al. (2024).
Knowledge Module. Some geometry problem solving systems integrate knowledge modules based on deep neural networks to more efficiently retrieve and apply knowledge and theorems in the field of geometry and verify the correctness of solutions. Knowledge Extractor and Integrator extracts and integrates geometric knowledge into the solution process. Some studies build geometric knowledge frameworks by constructing knowledge graphs. Fu et al. (2019) and Zhou et al. (2022) use BiLSTM to extract geometric relationships, and Tsai et al. (2021) uses Graph Convolutional Network (GCN) Kipf and Welling (2017) to embed knowledge graphs into vector space.  Xu et al. (2024b) and  Sharma et al. (2025) use CLIP and VISTA Zhou et al. (2024a) models to encode geometry problems to retrieve similar problems. Xiao et al. (2024a) builds a complete knowledge system through LSTM. Theorem Predictor is used to predict the geometric theorems required for the current solution step to derive a formal solution path. Guo and Jian (2022); Jian et al. (2023a) encodes the structural information of the formal language through GCN and uses the Sequence-to-Sequence (Seq2Seq) architecture Sutskever et al. (2014) of BiLSTM-GRU for theorem prediction. In addition, many studies use the Transformer-based Seq2Seq architecture for prediction Lu et al. (2021); Wu et al. (2024a); Zhang et al. (2024h), and some of these introduce the T5 model Yang et al. (2023); He et al. (2024b); Shengyuan and Xiuqin (2024). Zou et al. (2024) leverages DistilBERT Sanh (2019) to guide the training of theorem predictors. Answer Verifier is used to ensure the correctness of the solution logic. In addition to the traditional rule-based verification method Zhang et al. (2024f), Pan et al. (2025) introduces a pre-trained large language model Qwen et al. (2025) to verify the solution steps.
3.1.2 Other Architectures
In addition to the encoder-decoder architecture, some deep learning systems for solving geometry problems have adopted other architectural forms. Song et al. (2023) combines a GAN with the counter-factual evolutionary reasoning mechanism, and Peng et al. (2023); Huang et al. (2024) use the GNN framework to solve geometry problems. Many studies adopt Decoder-Only Architecture to solve unimodal geometry problems. The AlphaGeometry series Trinh et al. (2024); Sinha et al. ; Chervonyi et al. (2025) uses a trained Transformer to solve geometry theorem problems at the IMO level. There are also studies that directly use LLMs such as Llama 2 Touvron et al. (2023), Mistral Jiang et al. (2023), and Code Llama Roziere et al. (2023) to solve geometry problems Cai et al. (2024a); Tong et al. (2024); Tang et al. (2024b); Li et al. (2024d).
Other studies have combined LLMs with other deep learning architectures, or multiple LLMs, to build Hybrid Architectures to solve geometry problems. Pi-GPS Zhao et al. (2025a) uses PGDPNet Zhang et al. (2022b) and MLLM to disambiguate diagrams and texts, and further uses o3-mini OpenAI (2025) to complete theorem prediction and problem solving; GeoUni Cheng et al. (2025a) uses DeepSeek-R1-Distill-Qwen Guo et al. (2025) combined with CNN encoders and decoders to generate and solve multimodal geometry problems. In addition, some studies adopt the multi-agent system consisting of multiple LLMs Zhang et al. (2024b); Gao et al. (2024); Lei et al. (2024); Liu et al. (2025b).
3.2 Training Stage for Geometry Problem Solving
3.2.1 Pre-Training
Pre-Training Task. In addition to directly using pre-trained models to solve geometry problems, many studies have designed targeted pre-training tasks to improve the performance of deep learning systems when dealing with geometry problems. Some pre-training tasks focus on the textual modality, Chen et al. (2022) proposes Mathematical Expression Pretraining (MEP) to learn mathematical knowledge, while Zhang et al. (2023a, 2024f); Ma et al. (2024a) use the Masked Language Modeling (MLM) task to enable the model to understand and generate textual descriptions of geometry problems effectively. There are also pre-training tasks that focus on diagram encoders. Chen et al. (2021) proposes Jigsaw Location Prediction (JLP) and Geometry Elements Prediction (GEP) tasks to improve performance, while Ning et al. (2023) optimizes diagram encoders through Masked Image Modeling (MIM) and Multi-Label Classification (MLC). Others focus on matching multimodal relationships. LANS Li et al. (2024h) uses two pre-training strategies, Structural-Semantic Pretraining (SSP) and Point-Match Pretraining (PMP), to achieve global relationship modeling and alignment between visual and text points. SANS Lin et al. (2025) uses Dual-Branch Visual-Textual Points Matching (DB-VTPM) to match points in vision and text, enabling the model to understand the spatial relationship between geometric primitives in geometric figures.
Pre-Training Data. Many studies have proposed solutions to the lack of geometric pre-training data. For example, AMPS Hendrycks et al. (2021) and InfiMM-WebMath-40B Han et al. (2024) are large-scale mathematical and multimodal pre-training datasets, respectively, which can significantly improve the performance of models on geometry tasks. In addition, existing pre-trained visual encoders and MLLMs are mainly trained based on real-world images, which are quite different from geometric diagrams Hsu et al. (2022); Sharma et al. (2025), so other studies have built special datasets for pre-training of diagram encoders. Geo-ViT Xia et al. (2025) collects more than 120K geometric diagrams for training on ViT; CLIP-Math Zhang et al. (2025c), GeoCLIP Cho et al. (2025a), GeoGLIP Zhang et al. (2025d), and the diagram encoder in DFE-GPS Zhang et al. (2025e) use large-scale training data synthesized by programs, specifically for geometric pre-training.
3.2.2 Supervised Fine-tuning
In geometry problem solving, deep learning models usually require Supervised Fine-Tuning (SFT) to achieve better performance, and the most critical factor is training data. In addition to collecting training data from textbooks, exams, and the Internet Seo et al. (2015); Lu et al. (2021), many studies have also carried out related work on the generation, augmentation, and filtering of training data.
Data Generation. Among the rule-based methods, Kim and Chun (2022); Trinh et al. (2024); Kamoi et al. (2024); Huang et al. (2025b); Zhuang et al. (2025) synthesizes data through a rule-based geometry problem generator, while Kazemi et al. (2024a); Zhang et al. (2025c); Sun et al. (2025c) use carefully designed program templates to construct complex geometric diagrams and corresponding problems from basic geometric elements. Recent work uses a multi-component pipeline to generate high-quality geometry question-answer pairs that include reasoning processes Pan et al. (2025); Fu et al. (2025). Among the LLM-based methods, Tang et al. (2024b); Huang et al. (2025e) generate new questions through mathematical knowledge points, and GeoUni Cheng et al. (2025a) is a geometry question generation framework based on LLM. Deng et al. (2024) uses a rule-based approach to generate geometry images and further uses an LLM to generate question-answer pairs. There are also research works using agents to generate new geometry problems Lee et al. (2025); Wen et al. (2025), among which the proof questions proposed by Tonggeometry Zhang et al. (2024b) were even selected for a competition.
Data Augmentation. Many studies have improved training results by enhancing existing datasets. For example, Cao and Xiao (2022); Zhang et al. (2023a, 2024d); Xiao and Zhang (2023); Lin et al. (2025); Zhuang et al. (2025) use rule-based data augmentation methods to diversify text and images, thereby improving the robustness of the model. Zhang et al. (2023c); Wu et al. (2024a) expand new questions through geometry theorems, while Tong et al. (2024); Shi et al. (2024); Anand et al. (2024a); Jaiswal et al. (2024); Cheng et al. (2024) use LLMs to generate diverse question-answer pairs of different difficulty levels. Some studies improve the reasoning ability of the model by introducing fine-grained annotated reasoning processes in the training data Cai et al. (2024a), including Chain-of-Thought (CoT) Gao et al. (2025b); Chen et al. (2024c); Sun et al. (2025b); Luo et al. (2025); Huang et al. (2025d); Ning et al. (2025), Program-of-Thought (PoT) Li et al. (2024d); Sharma et al. (2025) and long CoT Xu et al. (2024a); Xiang et al. (2024, 2025); Xu et al. (2025a); Du et al. (2025). Zhao et al. (2024); Cai et al. (2024b) enhance the geometric understanding of models by drawing geometric diagrams for unimodal problems or generating diagrams that better match the text. In addition, adding diagram descriptions (including diagram literals Tey and captions Gao et al. (2025b); Zhang et al. (2025e); Xia et al. (2025); Huang et al. (2025f)) to the training set can significantly improve the ability to understand geometry problems, especially enhancing the model’s attention to diagrams during reasoning.
Data Filtering. Sun et al. (2025b); Fu et al. (2025); Wang et al. (2025e) use search algorithms to screen data quality and difficulty, while Cai et al. (2024b); Han et al. (2024); Luo et al. (2025); Jia et al. (2025); Huang et al. (2025e) use LLMs to score samples to screen out high-quality data.
3.2.3 Reinforcement Learning
Reinforcement Learning (RL) can significantly improve the geometric reasoning capabilities of deep learning models.
Non-LLM Algorithms. Some studies have used Deep Reinforcement Learning (DRL) methods without LLM to solve geometry problems Zou et al. (2024), such as the Deep Q-Network (DQN) Mnih et al. (2013) algorithm Peng et al. (2023); Huang et al. (2024) and the Proximal Policy Optimization (PPO) Schulman et al. (2017) algorithm Xiao and Zhang (2023).
LLM Algorithms. Many works using LLM usually introduce RL after SFT. Common algorithms include PPO Peng et al. (2024, 2025), Direct Preference Optimization (DPO) Rafailov et al. (2023); Zhang et al. (2025c); Xu et al. (2025a); Huang et al. (2025b), and Group Relative Policy Optimization (GRPO) Guo et al. (2025); Huang et al. (2025d); Deng et al. (2025b); Tan et al. (2025); Deng et al. (2025a); Cheng et al. (2025a); Huang et al. (2025c); Chen et al. (2025b); Liu et al. (2025a); Wang et al. (2025i). In addition, some studies have adopted other algorithms such as Group Policy Gradient (GPG) Chu et al. (2025).
3.3 Inference Stage for Geometry Problem Solving
3.3.1 Test-Time Scaling
In the inference stage, Test-Time Scaling (TTS) has become a research area that has attracted much attention recently. It is widely used to significantly improve the reasoning ability of the model.
X-of-Thought. Various X-of-Thought methods essentially encourage LLMs to generate longer and more diverse samples to guide the model to invest more computing resources, rather than just generating a single short sample like traditional methods Zhang et al. (2025b). Many studies have adopted different CoT Wei et al. (2022) methods to solve geometry problems Xu et al. (2024c); Fu et al. (2024); Taveekitworachai et al. (2024), some of which involve multiple rounds of interaction with the model Zhang et al. (2023d); Zheng et al. . To improve arithmetic accuracy, some studies have introduced PoT Chen et al. (2023a) methods, including directly generating complete programs DAS et al. (2024); Chen et al. (2024b) or generating multiple subprograms through distribution Singh et al. (2025). In addition, Duan et al. (2024); Liu et al. (2023) have combined multiple X-of-Thought methods such as CoT and PoT, and some have solved problems by combining CoT with external tools Qian et al. (2023); Gou et al. (2024). There are also some multimodal CoT methods, which generate formal language descriptions of geometric diagrams Zhou et al. (2024c) or natural language descriptions Jia et al. (2024); Tey ; Singh et al. (2024) first, and then perform reasoning.
Search Methods. Many works combine deep learning systems for geometry problem solving with a variety of search algorithms to achieve more robust reasoning. These works mainly use various tree search algorithms Zhang et al. (2024b); Pan et al. (2025), such as Beam Search Trinh et al. (2024); Chervonyi et al. (2025); Peng et al. (2023); Zhang et al. (2024d); Xu et al. (2024a), Monte Carlo Tree Search (MCTS) Coulom (2006); Zou et al. (2024); Rabby et al. (2024); Yao et al. (2024); Dong et al. (2024); Wu et al. (2025), and Predictive Rollout Search (PRS) Wang et al. (2025f), to improve reasoning performance significantly. In addition, Xiong et al. (2024) uses a graph search algorithm for reasoning.
Verification Methods. A reliable verification method is crucial in test-time scaling. Related studies usually use the Process Reward Model (PRM) to evaluate the quality of the reasoning steps Xiang et al. (2024, 2025); Luo et al. (2025); Wang et al. (2025c); Tu et al. (2025); Dong et al. (2024); Hu et al. (2025), and often combine it with the search algorithm as a provider of path confidence. In addition, Yuxuan et al. (2024) uses a method based on generation logits to measure the confidence of the reasoning steps, while Zhang et al. (2025a) trains an outcome verifier to evaluate the accuracy of the reasoning results.
Others. Cheng et al. (2024) uses an LLM to filter out the correct answer from multiple generated candidate solutions, while Jin et al. (2025) proposes an agent to dynamically manage the number of multi-agents, their reasoning strategies, and how they collaborate.
3.3.2 Knowledge-Augmented Inference
Knowledge-augmented inference aims to enhance the reasoning performance of models by incorporating external sources of knowledge.
Few-shot Learning. Few-shot Learning Brown et al. (2020) is also an important prompting method that can guide the model to solve similar geometry problems. Several studies provide examples through In-Context Learning (ICL) Agrawal et al. (2024); Cheng et al. (2025b), some of which provide examples from the perspective of basic skills Chen et al. (2024a), some incorporate curriculum learning methods Vu et al. (2025), and some place text in images Wang et al. (2024b). In addition, some studies adopted the Retrieval-Augmented Generation (RAG) paradigm to retrieve similar examples through external databases to provide hints Xu et al. (2024b); Jaiswal et al. (2024); Sharma et al. (2025).
Visual Aids. For geometry problems, some studies start with the corresponding geometric diagrams and process the diagrams during the inference process to help solve the problem. Xiao et al. (2024b) uses drawing tools to convert text problems into dual-modal representations of vision and language for reasoning, while Hu et al. ; Chen et al. (2025c); Qi et al. (2025); Wang et al. (2025f) assist in solving problems by drawing auxiliary lines on geometric diagrams or highlighting key features.
Others. Guo et al. (2024b) uses learned task plans to guide reasoning, while Yin et al. (2024) solves geometry problems by updating explicit memory, leveraging important contextual knowledge captured during training.
4 Evaluations for Geometry Problem Solving
In this section, we summarize the evaluation metrics and methods for geometry problem solving.
4.1 Evaluation Metrics
Evaluation metrics include performance-based metrics (outcome-based metrics, process-based metrics), and efficiency-based metrics.
4.1.1 Performance-Based Metrics
Outcome-Based Metrics. Outcome-based metrics focus on measuring the correctness of the final inference results without considering the details of the reasoning process. Top-k Accuracy (Top-k Acc) and Pass@n (P@n) are two main metrics for measuring answer accuracy. Top-k Acc refers to the ratio of the first k predictions generated by the model that have at least one correct answer, while P@n refers to the ratio of the model that can correctly solve the problem at least once in a given n attempts. Other works also use outcome-based metrics such as F1 score Mishra et al. (2022b); Cheng et al. (2025b), maj@k Yue et al. (2024a), number of correct and wrong answers Dou et al. (2024), choice Zhang et al. (2023a) and competition scores. F1 score is a metric that takes into account both precision and recall; maj@k measures the ratio of cases that the model gets the correct answer through majority vote in k samples; the number of correct and wrong answers intuitively shows the task results; choice refers to the ratio of correctly selected answers from four candidate options of a multiple-choice question, with a random selection made if the model’s output does not match any of the given options; some studies also show model performance by showing SAT scores Seo et al. (2015) or IMO competition scores Trinh et al. (2024).
Process-Based Metrics. Recently, as the demand for improving model performance continues to increase, more and more research has begun to focus on the reasoning process of deep learning systems, rather than just the final result. Completion Zhang et al. (2023a) and No Result Chen et al. (2021) measure the executableness of the reasoning process: Completion refers to the correct rate of selecting the first executable solution of the deep learning model as the result; No Result indicates the ratio of the inference program failing to produce a result. For questions that contain standard reference process answers (i.e., CoT), several studies have adopted a variety of reasoning process evaluation strategies to judge the quality and correctness of the reasoning process Jaiswal et al. (2024); Qiao et al. (2024), including N-gram Similarity Ma et al. (2024b), Step Accuracy Rate Wang et al. (2025b), and CoT-E score Chen et al. (2025a). For process-based metrics that are difficult to quantify, such as the precision of each reasoning step when there is no reference reasoning process Zhang et al. (2024g); Liu et al. (2024c); Zhou et al. (2024b), the logical coherence of CoT Zhang et al. (2025e), and the interpretability of the reasoning process Sachan et al. (2017); Sachan and Xing (2017), manual evaluation or LLM evaluation methods are usually used for scoring.
4.1.2 Efficiency-Based Metrics
Efficiency-based metrics are used to measure the resource consumption and efficiency performance of the model during reasoning. These metrics include the time it takes to solve the problem Alvin et al. (2017), the ratio of unsuccessful solutions within the specified time (timeoutZhang et al. (2023c), the number of inference steps Wu et al. (2024a); Fang et al. (2024), and the cost of running the model Balunović et al. (2025).
4.2 Evaluation Methods
Evaluation methods related to geometry problem solving can be divided into three categories: manual evaluation, rule-based evaluation, and evaluation using LLMs (LLM-as-a-judge).
4.2.1 Manual Evaluation
In geometry problem solving, manual evaluation is relatively rarely used. It is done by human experts or annotators who directly check the output or reasoning process of the model. The core uses of manual evaluation include: first, evaluating the correctness of complex answers (e.g., judging whether $\frac{1}{\sqrt{2}}$ is equivalent to $\frac{\sqrt{2}}{2}$u et al. (2023); second, evaluating the interpretability of the reasoning process Sachan et al. (2017); Trinh et al. (2024). In addition, many studies will manually check the reasons for wrong and correct answers, which is also called a case study Lu et al. (2021); He et al. (2024a).
4.2.2 Rule-Based Evaluation
Rule-based evaluation is widely adopted in geometry problem solving tasks. It automatically detects whether the output meets the standard through clearly designed rules or programs, and has the advantages of being efficient, objective and repeatable. Common methods include extracting answers using regular expressions Gao et al. (2025b), comparing equivalence expressions using a math parser Yue et al. (2024a), normalizing answers, and removing dimensions and units Sun et al. (2024). In addition, some other methods include using the running results of the Python code as the answer Kim et al. (2023), and allowing the numerical answer to have errors within a certain range Kazemi et al. (2024a). In order to improve the stability of the evaluation, some studies have adopted majority voting to select the answer with the highest probability, or enhanced the stability of the results by taking the average of multiple evaluations Yan et al. (2025).
4.2.3 LLM-as-a-Judge
Recently, as the capabilities of LLMs continue to increase, more and more studies have begun to use the “LLM-as-a-judge” paradigm for evaluation rather than rule-based methods Li et al. (2024a). This paradigm is not only more cost-effective than manual evaluation, but also more scalable than rule-based evaluation, and is suitable for complex tasks such as geometry problem solving. Common uses include extracting answers Lu et al. (2024); Zhang et al. (2024e), standardizing answers Gao et al. (2025a), evaluating the reasoning process Anand et al. (2024a); Jaiswal et al. (2024), and checking errors Hendrycks et al. (2021); Sun et al. (2024). When evaluating the quality of the reasoning process, because some questions may have multiple solutions, rigid standard answers may limit the accuracy of actual reasoning evaluation. Using LLM to extract the key steps of the reasoning process and score them one by one is more flexible and more in line with real scenarios Zhang et al. (2024g). However, we need to be wary of the possible hallucinations and extra resource consumption that LLM may cause.
5 Discussion
5.1 Challenges
Data. First, there are significant deficiencies in the current geometry problem solving data. From the perspective of task type, compared with geometric numerical calculation, there is a serious lack of data on geometry theorem proving. From the perspective of geometry type, compared with plane geometry, there is also a lack of data on solid geometry and analytic geometry. From the perspective of language type, the data is concentrated in English and Chinese, and there is almost no data in other languages. Second, there is still a large gap between synthetic data and real exam questions. Although some methods of synthesizing geometric data have been proposed recently to generate large-scale data for model training, the improvement in model performance after adding these synthetic data to the training set is still limited Pan et al. (2025); Fu et al. (2025). This suggests that we need to further think about how to synthesize more realistic and effective data to better help the model solve geometry problems in the real world. In addition, current datasets generally lack annotations for intermediate steps and reasoning processes Shi et al. (2024), which is also a direction for improvement in future work. More discussion of the data can be found in the Appendix A.
Evaluation. First, the question types are monotonous. Existing benchmarks mainly use multiple-choice questions for evaluation (see Table 1), which provides space for models to guess answers, thus affecting the accuracy and reliability of the evaluation results. Although some studies have attempted to alleviate this issue—such as by evaluating different permutations of answer choices to mitigate positional bias (as in CircularEval) Liu et al. (2024b), or by removing candidate options entirely Fu et al. (2025)—these strategies have not yet been widely adopted. Second, there is no standard method for evaluating the reasoning process. Recently, with the increasing demand for model improvement, the evaluation of the geometry reasoning process has gradually received attention. However, these evaluation methods lack unified standards, and we should formulate more precise evaluation standards to accurately identify and improve the deficiencies of the model Park et al. (2024). In addition, current benchmarks may lack robustness, and the performance of the model is often not stable enough after applying small perturbations to the same evaluation questions Wang et al. (2025d); Zhou et al. (2024d); Weng et al. (2024). At the same time, some datasets may be included in the training data, thus affecting the true evaluation of model performance Park et al. (2024). This requires the academic community to introduce a fairer and more authoritative evaluation method.
Capability. Current deep learning systems still have many deficiencies in their capabilities when it comes to solving geometry problems. Given that most geometry problems involve multimodal inputs, models’ geometric visual perception capabilities are particularly crucial. However, studies have shown that introducing geometric figures often decreases model accuracy compared to using text input alone Zhang et al. (2025e); Onuoha et al. (2025). In multimodal geometry problems, spatial perception of geometric diagrams remains a major bottleneck, severely limiting problem-solving performance Sun et al. (2024); Xing et al. (2024); Kamoi et al. (2024); Zhang et al. (2024c). Some studies point out that deep learning models struggle with detecting Okada et al. (2023); Cho et al. (2025a) and perceiving Wang et al. (2025d); Weng et al. (2025) geometric angles. Others suggest that models also lack accurate perception of line lengths Wei et al. (2024); Huang et al. (2025b). This may be due to the inherently one-dimensional, sequential nature of current model architectures, which limits their ability to grasp the kind of spatial perception humans intuitively possess Sun et al. (2025d). Another reason could be the low resolution of visual encoders Zhang et al. (2024a); Zhu et al. (2025), which are primarily trained on natural images Hsu et al. (2022); Sharma et al. (2025), leading to suboptimal performance on geometric figures. Furthermore, many deep learning models also struggle with arithmetic accuracy. Some models adopt symbolic or formal reasoning methods Ning et al. (2025); Chervonyi et al. (2025), while others incorporate external computation modules to execute code and mitigate this issue Duan et al. (2024); Zhang et al. (2024f); Pan et al. (2024). In addition, LLMs may also exhibit a certain mindset when solving geometry problems, such as a tendency to construct coordinate systems as a default strategy Sun et al. (2025d). However, this approach can fail when it does not fit the context of the specific problem.
5.2 Future Directions
Combining Perception and Reasoning. Studies pointed out that visual perception errors and reasoning errors are the two main reasons for model errors Park et al. (2024); Wang et al. (2025b). Many studies have improved the performance of multimodal geometry problem solving from the perspectives of perception or reasoning. Early studies mainly focused on enhancing problem-solving capabilities through reasoning strategies, while recent studies have gradually shifted their focus to improving perception capabilities. In the field of perception, researchers have developed visual encoders Zhang et al. (2025d); Cho et al. (2025a) and models Chen et al. (2025a) specifically for geometric diagram understanding; in the field of reasoning, many works have tried various X-of-Thought methods or adopted training strategies such as RL. However, how to organically combine perception and reasoning is still a problem that needs to be solved. In addition, Sun et al. (2025c) points out that the two are not mutually exclusive, but rather complementary. For example, better modality alignment tasks can be designed for specialized visual encoders or modules to stimulate the model’s reasoning ability; more efficient multimodal CoT methods can be explored to achieve a deep fusion of perception and reasoning; and more reasonable RL training strategies, training set design methods, and reward mechanisms can be considered based on algorithms such as GRPO. It should be emphasized that the training set designed for SFT may not be suitable for RL Chen et al. (2025b), and this issue needs to be considered comprehensively from multiple perspectives such as diversity and generalization.
Using Cognitive Pattern. Cognitive pattern is a comprehensive approach that aims to simulate the cognitive process of humans in understanding and solving complex problems Kurbatov et al. (2021, 2022). The concept appeared in early problem-solving research. In the study of geometry problem solving, many strategies that simulate the human problem-solving process have been proven to be effective Zhou and Yu (2021); Rao et al. (2022). These strategies include: highlighting key information in diagrams and texts to focus on the key points of the problem; referencing diagram annotations, or adding auxiliary lines, coordinate axes and other elements to the diagram to enhance the understanding of the geometric structure; searching and applying relevant geometry theorems and knowledge to guide the problem-solving process; and using curriculum learning methods to improve the model’s problem-solving ability gradually. These methods have not yet been fully integrated into current advanced deep learning systems and deserve further attention and exploration.
Educational System. Before deep learning was widely used, there were already many systems and tools for geometry education, such as automatic scoring systems Mendis et al. (2017), automatic theorem discovery systems Kovács and Yu (2021), automatic problem solving systems Kang et al. (2016); Kurbatov et al. (2020); Kurbatov (2021); Kurbatov and Fominykh (2022); Li et al. (2024c), etc. These tools are designed to help teachers and students better learn and teach geometry knowledge. However, in the era of deep learning, systems or intelligent agents for geometry education are relatively rare. Automated geometry problem solving is considered one of the important development directions of intelligent education in the future Yang et al. (2023). Current artificial intelligence tools have made significant progress in solving geometry problems, but they still face many challenges in becoming easy-to-use teaching tools, including multi-language support and visual interaction. The capabilities of these tools are still relatively limited when dealing with geometry problems in the real world. In addition, intelligent agents specifically for educational scenarios are still relatively scarce, and more research work is urgently needed to address the complex challenges in the field of education.
6 Conclusion
In this paper, we provide a comprehensive and systematic survey of geometry problem solving. We summarize the relevant tasks, deep learning architectures and other methods, evaluation metrics and methods in detail, and provide an in-depth analysis of the shortcomings of existing data, evaluation, and model capabilities. Finally, we look forward to possible future research directions and emphasize the vast space for exploration in this field. This article aims to provide readers who are interested in this field with a comprehensive and practical resource to meet their research needs.
Limitations
Our survey focuses on the intersection of deep learning and geometry problem solving tasks in the past decade, and may not be able to fully present the development process of the entire field. In addition, given the rapid development of this field, our survey may not be able to timely reflect the latest developments and progress before and after the survey. Furthermore, our survey is mainly dedicated to summarizing existing research work, and there are deficiencies in experimental analysis. Despite these limitations, this survey still provides a valuable overview of the current status and main trends in the field of deep learning for geometry problem solving, which is expected to provide a useful reference for researchers and practitioners in this field.
Appendix A Geometry Problem Solving Datasets
In this section, we will further analyze various datasets for geometry problem solving. Table 1 and Table 2 provide a comprehensive summary of the datasets for geometry problem solving tasks from multiple perspectives, including dataset name, task type, geometry type, grade, problem source, whether it contains pictures, language, question type, rationale, training set and validation set size, test set size, and whether it is open source (if the data is open source, click the check mark to jump to the corresponding link).
There is a lack of geometry theorem proving data. Current academic research is mainly focused on geometric numerical calculation, while there is relatively little research on geometry theorem proving, and related data resources are also relatively scarce. There are many similarities between these two in terms of problem statement and mathematical knowledge Chen et al. (2022), but proof questions and calculation questions each have their own characteristics, and we should give them equal attention.
There is a lack of solid geometry and analytic geometry data. The data currently used for geometry problem solving tasks is mainly concentrated in the field of plane geometry, while for other types of geometry problems, such as solid geometry Yu et al. (2021b) and analytic geometry Wu et al. (2023), there is still a lack of open-source training and testing data. A study points out that current solid geometry problems are often relatively simple and regular Xu et al. (2025b). Their diagrams usually only contain simple visual elements and rarely involve complex combinations of multiple geometric bodies, which to a certain extent limits the development of the field of solid geometry. Even in the field of plane geometry, high-quality data available for evaluation is relatively scarce recently.
There are limitations in data sources. Although these datasets are authentic and reliable, they are limited in quantity. Recently, due to the scarcity of real data and copyright protection issues, large-scale datasets are mostly constructed using data augmentation and program synthesis techniques Gao et al. (2025b); Pan et al. (2025). However, the data generated by these methods is insufficient in terms of authenticity, diversity, and quality, making it difficult to completely replace data from real scenarios.
There are deviations in language types and question types. In terms of language, existing data is mainly concentrated in English and Chinese, while real data for native speakers of other languages is extremely scarce Zhang et al. (2023b); Song et al. (2024), which not only limits the evaluation opportunities for relevant national exams, but also affects the fairness of the evaluation. In addition, the question types are mostly multiple-choice questions, which provide space for the model to guess the answer and are not conducive to the accurate evaluation of deep learning models.
The datasets are also deficient in rationale. Most datasets lack detailed annotations of intermediate steps and reasoning processes Shi et al. (2024). Even if rationales exist, their formats are not uniform and granular enough to meet the goal of evaluating reasoning steps. In addition, the annotation format is mostly in natural language, which may not meet the needs of deep learning systems that reason in formal languages.
Datasets Task Type Grade Source Image Language Question Rationale Trainval Size Test Size Opensource
Fundamental Tasks
GeoC50 (2017) RE P - exist (dataset) zh FR - - 50
2Dgeometricshapes (2020) ER P - program - CQ - 36000 54000
GeoRE (2021a) RE P 6-12 Internet zh FR - 10000 2901
PGDP5K (2022; 2022b; 2022a) DP P 6-12 exist, textbook en FR - 4000 1000
Geoclidean (2022) SR P - program en YN - 185 555
BBH-geometricshapes (2023) ER P - program en MC - - 250
Tangram (2024a) ER P, S 1-12 exam, textbook en NR - - 4320
GeoCQT (2024a) ER P - exist, textbook - CQ - $\sim$11200 $\sim$2800
SP-1 (2024) DP P - program en FR - 200000 480
ElementaryGeometryQA (2024) DP, SP P 1-5 textbook en FR - - $\sim$500
Geoperception (2024c) SR P 6-12 exist en SA - - 11657
GePBench (2024) ER, SR P - program en MC - $\sim$300000 285000
CurveML (2024) ER P - program - CQ - 468000 52000
AVSBench (2024) ER, SR P - program en MC, FR - - 5073
VisOnlyQA (2024) ER, SR P - exist, program en MC, YN - 70000 1600
AutoGeo-100k (2025f) DC P - program en FR - 100000 -
Geo170K-alignment (2025b) DC P 6-12 exist en FR - 60252 -
GeomRel (2025d) SR P - program en MC - - 2629
ElementaryCQT (2025) ER P - program - CQ - 342000 38000
SynthGeo228K (2025e) DP, DC P - program en FR - 205491 22833
formalgeo-structure774k (2025e) DP, DC P 6-12 exist en FR - $\sim$774000 -
VGPR (2025a) ER, SR P - program en MC - 300000 50000
GeoX-alignment (2025) DC P - Internet en FR - 6232 -
VisNumBench (2025) ER, SR P - exist, prog, web en MC - - 1913
GeoPeP (2025c) ER, SR P, S - program en FR nl 200000 -
MathGlance (2025c) ER, SR P, S - exist, program en MC, YN, FR - - 1609
CogAlign-Probing (2025b) SR P - program en YN - 44000 4000
CogAlign-train (2025b) SR P, S - program en FR - 64000 -
Core Tasks
GEOS (2015) NC P 6-10 exam en MC - 67 119
GeoShader (2017) NC P 6-10 textbook, exam en NR - - 102
GEOS++ (2017; 2019) NC P 6-10 textbook en MC - 500 906
GEOS-OS (2017) NC P 6-10 textbook en MC demonstration 2235 -
Geometry3K (2021) NC P 6-12 online library en MC - 2401 601
GeoQA (2021) NC P 6-12 exam zh MC program 4244 754
Geometry3Dcalculation (2021b) NC S - website en, zh NR - - 140
Proving2H (2021) TP P 6-9 textbook, Internet zh FR - - 110
GeometryQA (2021) NC P 1-6 exist zh NR equations 1118 280
GeoQA+ (2022) NC P 6-12 website zh MC program 12054 -
UniGeo (2022) TP, NC P 9-12 website, exist en MC, FR program 12340 2201
BIG-bench-IG (2022) NC P - program en NR - - 250000
PGPS9K (2023a) NC P 6-12 exist, textbook en NR program 8022 1000
formalgeo7k (2024i) NC P 6-12 exist en, zh NR formal $\sim$5934 $\sim$1047
formalgeo-imo (2023c) TP P - online en, zh FR formal - 18
Conic10K (2023) NC A 10-12 website zh FR nl 8793 2068
GeomVerse (2024a) NC P - program en NR nl 11190 29000
IMO-AG-30 (2024) TP P - exam en FR - - 30
aug-Geo3K (2024a) NC P 6-12 exist en MC nl 13783 3824
GeoEval (2024e) NC P, S, A 1-12 exist, online en MC - - 5050
GeoGPT4V-GPS (2024b) TP, NC P 6-12 exist en, zh MC, FR nl 16557 -
GeoVQA (2024a) TP, NC P, S 6-12 textbook en NR, FR nl 4440 150
GeoMath (2024b) TP, NC S 10-12 website en NR, FB, FR nl 9155 906
GeoMM (2024) NC P - program en NR nl 87000 -
GPSM4K (2024b; 2024) TP, NC P 7-12 textbook en NR, FR nl 4272 1068
NBLP (2024) NC P 7-9 textbook, exam en NR, YN - - 100
G-MATH (2024) NC P, S 9-12 exist en FR - - 187
MathCheck-GEO (2024d) MR P 6-12 exist en NR, YN, FR nl - 1440
MO-TG-225 (2024b) TP P - exam en FR - - 225
Geo170K-qa (2025b) NC P 6-12 exist en MC nl 117205 -
FormalGeo7K-v2 (2025) NC P 6-12 exist en, zh NR formal 5950 1050
VerMulti-Geo (2025) NC P 6-12 exist en MC - 15000 -
GeoMath-8K (2025) NC P 6-12 exist en MC - 4500 820
GNS-260K (2025) KP, NC P 6-12 exist en MC, NR, SA program, nl 260017 -
GeoExpand (2025) TP, NC P 6-12 exist en MC, FR nl 45526 -
GeoSynth (2025) TP, NC P - program en MC, FR nl 62868 -
IMO-AG-50 (2025) TP P - exam en FR - - 50
GeoTrust (2025) NC P - program en NR nl $\sim$200000 240
GeoSense (2025b) KP, NC P, S 6-12 exist, website en, zh MC, FR - - 1789
formalgeo-reasoning238k (2025e) NC P 6-12 exist en NR nl $\sim$238000 -
Table 1: A summarization of geometry problem solving datasets for fundamental tasks and core tasks. Task: ER: Geometric Element Recognition, SR: Geometric Structure Recognition, DP: Geometric Diagram Parsing, DC: Geometric Diagram Captioning, SP: Semantic Parsing for geometry problem texts, RE: Geometric Relation Extraction, KP: Geometric Knowledge Prediction, TP: Geometry Theorem Proving, NC: Geometric Numerical Calculation. Type: P: plane geometry, S: solid geometry, A: analytic geometry. Question: MC: Multiple-Choice, NR: Numerical Response, FR: Free-Response, FB: Fill-in-the-Blank, YN: Yes-or-No, SA: Short-Answer, CQ: Classification Question. Rationale: nl: natural language. indicates that the dataset contains more than just geometry related content.
Datasets Task Type Grade Source Image Language Question Rationale Trainval Size Test Size Opensource
Composite Tasks
MATH (2021) MR P, S 9-12 exam en NR nl 7500 5000
AMPS (2021) MR P, S 1-12 website, program en NR, FR nl $\sim$5100k -
NumGLUE (2022b) MR P 6-10 exist en CQ - 81466 10583
Lila (2022a) MR P, S - exist en MC, FB, FR program 107052 26763
DMath (2023) MR P 1-6 handcraft en, kr NR program 7943 2079
TheoremQA (2023b) MR P - Internet, expert en MC, YN, FR - - $\sim$350
M3Exam (2023b) MR P, S 1-12 exam 9 lang. MC - - 12317
OlympiadBench (2024a) MR P, S - exam en, zh NR, FR nl - 8476
MathVista (2024) MR P 6-12 exist en, zh MC, FR - - 6141
MathVerse (2024g) MR P, S - exist, website en MC, FR nl - 15672
Math-Vision (2024a) MR P, S, A 1-12 exam en MC, FR - - 3040
MM-Math (2024) MR P 7-9 website en FR nl - 5929
We-Math (2024) MR P, S 3-6 website en MC - - 6524
VisAidMath (2024b) MR P, S, A 7-12 exam en MC, FR, YN - - 1200
CMM-Math (2024c) MR P, S, A 1-12 exam zh MC, FB, YN, FR nl 22248 5821
MathScape (2024b) MR P, S 1-12 homework, exam en NR, FB, FR nl - 1325
VisScience (2024) MR P, S 1-12 - en, zh MC, FR - - 3000
ArXivQA (2024b) MR P - paper en MC nl 100000 -
ReMI (2024b) MR P - - en MC, NR, FR - - 2600
MathV360K (2024) MR P 9-16 exist en MC, NR, FR - 360000 -
MultiMath-300K (2024) MR P, S 1-12 textbook, exam en, zh FB, NR, FR nl 290227 8443
InfiMM-WebMath-40B (2024) MR - - website en, zh - - $\sim$24000k -
MathVL (2024b) MR P, S, A 1-12 exist, private en MC, FB, FR nl 484914 2000
ArMATH (2024) MR P 1-6 school ar FR - - 200
M3CoT (2024c) MR P - exist en MC nl 9100 2359
MathOdyssey (2024) MR P, S 10-16 expert en MC, YN, FR nl - 387
PutnamBench (2024) MR - - exam en FR formal - 1692
ConcepetMath (2024b) MR P, S 1-9 website, textbook en, zh NR - - 4011
MATH() (2024) MR P, S 9-12 exist en NR - - 2060
MathBench (2024b) MR P, S 1-16 exist, exam en, zh MC, FR - - 3709
HARP (2024a) MR P - website en MC, SA, FR nl - 5409
M3GIA (2024) MR P 6-12 exam 6 lang. MC - - 1800
DART-Math (2024) MR P, S 9-12 exist en NR nl $\sim$1180k -
MathScaleQA (2024b) MR P, S 1-16 exist, exam en FR nl 2000000 -
UTMath (2024a) MR P, S - OEIS en NR - - 1053
MultiLingPoT (2024d) MR P, S 9-12 exist program NR program 41134 -
EITMath (2024a) MR P, S 9-12 exist en NR nl 15000 -
AIME2024 (2024) MR P, S - exam en NR nl - 30
AMATH-SFT (2024; 2025) MR P - exist en MC, FR nl $\sim$124000 -
MMathCoT-1M (2025) MR P - exist en MC, NR, FR nl $\sim$1020k -
DynaMath (2025) MR P, S, A 1-16 exist, website en MC, FR nl - 5010
CoMT (2025b) MR P - exist en MC nl - 3853
Diagramma (2025) MR P - program en MC - - 1058
MV-MATH (2025b) MR P, S, A 1-12 textbook, exam en MC, FR nl - 2009
CMMaTH (2025) MR P, S, A 1-12 website zh MC, FR nl - 23856
Math-PUMA-1M (2025) MR P, S - exist, online, prog en FR nl 996000 -
VisualWebInstruct (2025) MR P 1-16 exist, Internet en - nl 906160 -
MAVIS-Instruct (2025c) MR P, S, A - exist, program en MC, FR nl 834000 -
FlowVerse (2025a) MR P, S 9-12 website en, zh MC, FR nl - 2000
Omni-Math (2025a) MR P, S, D - exam en NR, FR nl - 4428
MathConstruct (2025) MR p 10-16 exam en FR - - 126
VCBench (2025j) MR P, S 1-6 textbook en MC - - 1720
OlymMATH (2025a) MR P, S, A - textbook, exam en, zh NR - - 200
RoR-Bench (2025) MR P, S 1-6 Internet zh FR nl - 215
PolyMath (2025g) MR P, S, A 1-16 exist, Internet 18 lang. NR - - 9000
MaTT (2025) MR P, S, A - reference book en MC - - 1958
CapaBench (2025) MR P, S 9-12 exist en NR nl - 1545
MATH-Perturb (2025a) MR P 9-12 exist en NR - - 558
M500 (2025) MR P - exam, exist en NR, FR nl 500 -
KPMATH-M (2025e) MR P, S 9-12 exist en NR nl 252000 -
Table 2: A summarization of geometry problem solving datasets for composite tasks. Task: MR: Mathematical Reasoning. Type: P: plane geometry, S: solid geometry, A: analytic geometry, D: differential geometry. Question: MC: Multiple-Choice, NR: Numerical Response, FR: Free-Response, FB: Fill-in-the-Blank, YN: Yes-or-No, SA: Short-Answer, CQ: Classification Question. Rationale: nl: natural language.
Appendix B Other Geometry Tasks
In addition to geometry problem solving, there are other geometry-related tasks that have not been systematically summarized, which have similar fundamental tasks. More details for the datasets can be found in Table 3.
Datasets Task Type Grade Source Image Language Question Rationale Trainval Size Test Size Opensource
Other Geometry Tasks
GMBL (2021) TD P - exam en GD - - 39
LeanEuclid (2024) AF P - exist, textbook en FR - 140 33
Euclidea (2024) CP P - website en FR nl - 98
PyEuclidea (2024) CP P - website program FR - - 98
MagicGeoBench (2025a) TD P 6-12 exam en GD - - 220
GeoX-pretrain (2025) DG P - web, textbook - GD - 127912 -
Table 3: A summarization of datasets for other geometry tasks. Task: TD: Geometric Text-to-Diagram; CP: Geometric Construction Problem; DG: Geometric Diagram Generation; AF: Geometric AutoFormalization. Type: P: plane geometry. Question: FR: Free-Response, GD: Geometric Diagram. Rationale: nl: natural language.
B.1 Geometric Diagram Generation
This task is dedicated to generating high-quality geometric diagrams, aiming to facilitate a deeper understanding of geometry problems and related applications such as image editing, thereby providing strong support for the field of education.
Geometric Diagram Reconstruction. This task is one of the earlier works in the field of geometry. It aims to use existing simple sketches or preliminarily drawn images to reconstruct a clearer and more standardized complete image, thereby helping users to understand and visualize the image content more intuitively Yu et al. (2015). One of the key challenges is to reconstruct 3D geometry from 2D single line drawing images Xue et al. (2010, 2012); Yang et al. (2013), even if the input image is incomplete or inaccurate Zheng et al. (2015, 2016b, 2016a); Zou et al. (2016).
Geometric Text-to-Diagram. This task requires the system to be able to generate corresponding geometric diagrams from the natural language description of the geometry problem. This ability will greatly enhance the understanding of the solution system, enabling it to more accurately grasp the flexible and changeable geometric propositions Liu et al. (2012). In addition to traditional rule-based methods Janičić and Narboux (2021); Krueger et al. (2021); Trinh et al. (2024), some recent studies have begun to use deep learning technology to build related systems Zhengyu and Xiuqin (2023); Wang et al. (2025a); Cheng et al. (2025a). MagicGeoBench Wang et al. (2025a) collects a dataset containing 220 plane geometry problems in middle school mathematics exams, specifically for evaluating the performance of text-to-diagram geometry generation models.
In addition to the above methods, there are other techniques for generating geometric diagrams. Some tools support generating drawings through interactive ruler and compass construction, including GeoGebra [^1] https://www.geogebra.org and Geometer’s Sketchpad Scher (1999). Besides, the are some non-interactive methods for deriving such constructions Bertot et al. (2004); Itzhaky et al. (2013). In order to support more forms of geometric diagram generation, some studies have tried a wider range of methods to construct geometric diagrams, including techniques such as algebraic numerical optimization Gao and Lin (2004) and constrained numerical optimization Ye et al. (2020).
This task is also related to geometry problem solving. GeoX Xia et al. (2025) builds a pre-trained dataset with more than 120,000 plane geometry images, and tunes the visual encoder-decoder architecture through the mask auto-encoding scheme to obtain a visual encoder that fully understands geometric images. Some geometry problem solving work uses related methods to perform data enhancement on single-modal geometry problems and generate corresponding diagrams to obtain multimodal data Cai et al. (2024b); Zhao et al. (2024); Xiao et al. (2024b).
B.2 Geometric Construction Problem
Geometric construction tasks, like geometry problem solving, are also part of educational exams. The goal is to use traditional ruler and compass construction methods to find an effective path to construct the target figure.
In recent years, some studies have tried to use deep learning systems to solve geometric construction tasks. In the online geometric construction game Euclidea [^2] https://www.euclidea.xyz,  Macke et al. (2021); Wong et al. (2023) uses Mask R-CNN He et al. (2017) to solve difficult geometric construction problems through a purely image-based method.  Mouselinos et al. (2024) converts the Euclidea problem into Python format and solves it with the help of a multi-agent framework based on large language models. This provides us with new ideas and inspires us to further explore the application potential of deep solution systems in cognitive fields such as planning and auxiliary line addition.
B.3 Geometric Diagram Retrieval
Before the widespread application of deep learning methods, the search problem of plane geometry diagrams has always been an important topic in the field of scientific research Liu et al. (2014a, b); Gan et al. (2016); Chen et al. (2016); Qu et al. (2016); Liu et al. (2016). With the development of computer technology, plane geometry retrieval may no longer be difficult in the era of deep learning. However, the retrieval of more complex solid geometry and irregular geometric figures may still be a direction worth studying.
B.4 Geometric Autoformalization
Autoformalization is a subtask of theorem proving Li et al. (2024f). A small number of studies focus on automatically converting informal geometry problems and proofs into formal theorems and proofs that can be verified by machines. LeanEuclid Murphy et al. (2024) is a 173-problem geometric autoformalization dataset designed to test whether AI can understand mathematical problems and solutions written by humans and convert them into formal theorems and proofs.
Paper Task Network Text Encoder Diagram Encoder Fusion Module Decoder Knowledge Module
Fundamental Tasks
RSP (2018) SP BiLSTM BiLSTM - - - -
2StepMemory (2020) SP attention attention - - - -
GIRTOOLS (2020) ER VGG16 - - - - -
Arsenal (2021) SP Seq2Seq RNN - - - -
PGDPNet (2022b) DP FPN+GNN - FPN+GNN - - -
UV-S2 (2022) RE - BERT RetinaNet - - -
BiLSTM-CRF (2022) RE BiLSTM - - - - -
Stacked LSTM (2022) DP LSTM - LSTM - - -
2DGeoShapeNet (2024b) ER CNN - - - - -
Euclid (2024c) SR MLLM - ConvNeXt MLP Qwen-2.5 -
FGeo-Parser (2025) DP, SP - T5 BLIP - - -
Core Tasks - Encoder-Decoder Architecture
Inter-GPS (2021) NC - - RetinaNet - - Transformer
NGS (2021) NC - LSTM ResNet101 co-attention LSTM -
S2G (2021) NC - BiGRU - - GRU GCN
GCN-FL (2022) NC - - DenseNet121+FPN - - GCN+BiLSTM-GRU
DPE-NGS (2022) NC - Bi-LSTM+RoBERTa ResNet101 co-attention LSTM -
Geoformer (2022) TP, NC MLLM - VL-T5 -
MCL (2023b) NC - BERT Faster R-CNN attention attention+GRU -
PGPSNet (2023a) NC - Transformer CNN BiGRU GRU -
UniMath (2023) TP, NC MLLM - VQ-VAE T5 -
RetinaNet+GCN (2023a) NC - - DenseNet121+FPN - - GCN+BiLSTM-GRU
SCA-GPS (2023) NC - Bi-LSTM+RoBERTa ViT co-attention LSTM -
TD-Parsing (2023) NC - - DenseNet121 co-attention LSTM -
SUFFI-GPSC (2023) NC - - - - - T5
LANS (2024h) NC - ResNet10+Transformer BiGRU GRU -
GAPS (2024d) TP, NC - - ResNet VL-T5 GRU -
E-GPS (2024a) NC - - PGDPNet - - Transformer
FGeo-TP (2024b) NC - - - - - Transformer
FGeo-DRL (2024) NC - - - - - DistilBERT
FGeo-HyperGNet (2024h) NC - - - - - Transformer
GOLD (2024) TP, NC MLLM - FPN+MobileNetV2+CNN T5 -
DualGeoSolver (2024a) NC - Bi-LSTM+RoBERTa ViTMAE co-attention LSTM LSTM
Math-LLaVA (2024) NC MLLM - LLaVA-1.5 -
PGPSNet-v2 (2024f) NC - Transformer+BiGRU CNN - GRU -
EAGLE (2024g) NC MLLM - LLaVA-1.5 -
MultiMath (2024) NC MLLM - CLIP-ViT MLP DeepSeekMath-RL -
MathGLM-Vision (2024b) NC MLLM - GLM-4V -
ATB-NGS (2024a) NC RoBERTa+BiLSTM Real-ESRGAN+ResNet101 co-attention LSTM -
Geo-Qwen (2024) NC MLLM - PGDPNet - Qwen2.5 T5
Geo-LLaVA (2024b) NC MLLM - LLaVA-1.5 CLIP
SANS (2025) NC - CNN+Transformer BiGRU GRU -
GNS-DTIF (2024a) NC - Transformer DenseNet121+GCN GRU+co-attention LSTM -
MATHS (2024) TP, NC - - Swin-Transformer - BERT -
G-LLaVA (2025b) NC MLLM - LLaVA-1.5 -
R-CoT (2024) NC MLLM - InternVL2 -
MAVIS (2025c) NC MLLM - CLIP-Math MLP MAmmoTH2 -
GeoX (2025) TP, NC MLLM - Geo-ViT GS-Former Geo-LLM -
DFE-GPS (2025e) NC MLLM - SigLIP MLP Yi-1.5 -
GeoDANO (2025a) NC MLLM - GeoCLIP MLP LLama-3 -
Math-PUMA (2025) NC MLLM - SigLIP MLP Qwen2 -
GeoCoder (2025) NC MLLM - LLaVA-1.5 VISTA
MAmmoTH-VL2 (2025) NC MLLM - MAmmoTH-VL -
GNS (2025) NC MLLM - LLaVA-1.5 -
GeoGen (2025) NC MLLM - Qwen2.5-VL Qwen2.5
RedStar-Geo (2025a) NC MLLM - InternVL2 -
SVE-Math (2025d) NC MLLM - GeoGLIP MLP Qwen2.5Math -
MGT-Geo (2025) NC MLLM - Qwen2.5-VL -
Core Tasks - Other Architecture
GAN+CfER (2023) NC cGAN - - - - -
GeoDRL (2023) NC GNN - - - - -
HGR (2024) NC GNN - - - - -
Table 4: A summarization of deep learning architecture for geometry problem solving system. Task: DP: Geometric Diagram Parsing, SP: Semantic Parsing for geometry problem text, ER: Geometric Element Recognition, DP: Geometric Diagram Parsing, SR: Geometric Structure Recognition, RE: Geometric Relation Extraction, TP: Geometry Theorem Proving, NC: Geometric Numerical Calculation. indicates the presence of the attention mechanism.